OpenStudy (anonymous):

Simplify. 2x(-11x - 7) A) –22x2 -14x B) –11x2 -14x C) -22x - 14x2 D) 22x2 -14x2

OpenStudy (anonymous):

pl rewrite the eq in the equation creator

OpenStudy (anonymous):

3/(2x) + x/(2x - 6) = 3/(2x) + x[2(x - 3)] = [3(x - 3) + x(x)]/[2x(x - 3)] = (3x - 9 + x^2)/[2x(x - 3)] = (x^2 + 3x - 9)/[2x(x - 3)